OpenStudy (anonymous):

what is the indefinite integral of: e^2x/(e^x + 1)^1/4

OpenStudy (anonymous):

Substitute \(u=e^x\) recognizing \(e^{2x}=(e^x)^2\).

jhonyy9 (jhonyy9):

\[e ^{(2x-(x+1)/4)}=e ^{((7x+1)/4)}\]